### v1

# CORT: Class-Oriented Real-time Tracking for Embedded Systems

Abstract
The ever-increasing use of artificial intelligence in autonomous systems has significantly contributed to advance the research on multi-object tracking, adopted in several real-time applications (e.g., autonomous driving, surveillance drones, robotics) to localize and follow the trajectory of multiple objects moving in front of a camera. Current tracking algorithms can be divided into two main categories: some approaches introduce complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance, whereas other approaches are aimed at reducing response times by removing the re-identification phase, thus penalizing the tracking accuracy. This work proposes a new approach to multi-class object tracking that allows achieving smaller and more predictable execution times, without penalizing the tracking performance. The idea is to reduce the problem of matching predictions with detections into smaller sub-problems by splitting the Hungarian matrix by class and invoking the second re-identification stage only when strictly necessary for a smaller number of elements. The proposed solution was evaluated in complex urban scenarios with several objects of different types (as cars, trucks, bikes, and pedestrians), showing the effectiveness of the multi-class approach with respect to state of the art trackers.
Index Terms:
Real-time tracking, multiple-object tracking, autonomous driving, embedded systems, real-time computing.
Figure 1: Urban scenario with six objects belonging to three different classes (car, bus, motorbike). The left diagram shows the standard monolithic formulation of the matching problem used in DeepSORT-like tracking algorithms, while the right diagram illustrates the proposed parallel class-based approach. In complex scenarios, the proposed solution is more resource efficient, exhibits time and memory scalability, and can also exploit multi-core CPUs to parallelize the smaller association problems.
I Introduction
Object tracking is a computer vision task for locating objects in a video sequence and generating their trajectories, associating each object not only with a bounding box, but also with a unique identifier. It is used in many application fields, including surveillance, autonomous driving, robotics, video analysis, and augmented reality. Differently from object detectors, which only analyze single frames, trackers process video sequences to reconstruct the trajectories of the moving objects (tracklets), which are used to re-identify objects in consecutive video frames, even in the presence of occlusions or crossing trajectories.
Before the advent of deep learning, multi-object tracking (MOT) approaches were based on correlation filters, like MOSSE [1] and ASEF [2], which primarily rely on handcrafted features, data association techniques, and linear motion models [3]. Later, the great performance of convolutional neural networks has revolutionized MOT by enabling end-to-end learning-based solutions, improving feature representation and object detection accuracy.
The power of artificial intelligence applied to object tracking became evident in 2017, with the first deep learning extension of the SORT algorithm [4], called DeepSORT [5]. Since then, object tracking architectures evolved to integrate more advanced neural networks, essential to enhance the capabilities of object tracking methodologies. The integration of deep learning models also lead to the development of novel tracking architectures, like those using recurrent neural networks (RNNs) [6, 7, 8] and graph neural networks (GNNs) [9, 10, 11]. However, the exceptional results achieved by recent object detection models make the tracking-by-detection paradigm the most effective approach for multiple object tracking.
Tracking by detection has rapidly spread into various application domains, including embedded systems, showing its potential in advanced driving assistance systems (ADAS) [12] and unmanned aerial vehicles (UAV) [13]. However, the integration of this technology into embedded devices faces a major issue due to the limited computational resources available in such systems. Most of the state-of-the-art approaches [5, 14, 15] focused on maximizing the tracking performance on common benchmarks, like the MOT-Challenge [16], introducing complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance. Other approaches [17, 18] reduced the response time of the tracking pipeline by removing the neural network for the re-identification phase, thus penalizing the tracking accuracy.
Contributions This paper presents CORT, a new solution for the multi-class MOT problem aimed at reducing execution times without penalizing tracking performance. In particular, this work provides the following novel contributions:
  • It proposes a parallel formulation of the matching problem by splitting the Hungarian matrix into smaller ones, using the object class as discriminant. As illustrated in Figure 1, this allows reducing the execution time of the overall matching phase to the one required to solve the largest sub-matrix. Due to the cubic complexity of the Hungarian matching algorithm, the proposed approach is particularly effective in the presence of objects of different classes.
  • It proposes a new approach for the cascade matching aimed at reducing execution times. In a first stage, a computationally efficient metric is used to solve the best-matching tracklets, while, in the second stage, the features extracted by a deep learning re-identification model are used for the remaining tracklets to handle long term occlusions and reduce ID-switches.
Paper structure The rest of the paper is organized as follows: Section II discusses the state of the art solutions related to this work; Section III presents the proposed approach; Section IV reports some experimental results aimed at showing the benefits of the proposed solution; and, finally, Section V states the conclusions and possible future developments.
II Related work
Multi-object tracking algorithms can be divided into two major categories: joint and separate trackers. Joint trackers integrate different components of the tracking pipeline in the same neural network, whereas separate trackers treat multiple objects independently, identifying objects in subsequent frames based on factors such as appearance, motion, and displacement.
II-A Joint trackers
Joint Detection and Embeddings (JDE) [19] proposes a neural architecture that combines object detection and visual appearance into a single shared model. However, directly fusing motion and appearance information in the cost matrix often leads to failures when the target is not detected for a few frames.
Tracktor [20] proposes a way to convert an object detector into a tracker by adding a regression head on top of the object detection model. It works by placing the bounding box information obtained at time $t-1$ in the current frame at time $t$ and extracting relevant features to refine the new object location, which becomes the actual tracked position. One problem of this method is that it works under the strong assumption that objects have minimal changes between consecutive frames. This means that either the objects should move at low speeds or the scene should be processed at high frame rates. However, a high frame rate is in contrast with the large computation time required for the inference of such a complex neural network. In fact, the complexity of the model limits the real-time capabilities of the algorithm. Processing frames at low rates is problematic. In fact, increasing the time between subsequent frames causes larger variations both in object positions and in their appearance. This makes the association step harder, possibly leading to a rapid performance degradation.
The problem of Tracktor was addressed in the same paper [20] by Tracktor++, which integrates a motion model for dealing with low frame rates and moving cameras, which is a typical issue in real-world scenarios. In spite of this extension, however, Tractor++ frequently fails in the presence of fast and small moving objects and is not suitable for real-time applications.
CenterTrack [21] proposes a different approach to the tracking problem. It is built on a modified version of the CenterNet [22] architecture. CenterNet generates a low-resolution heatmap, representing the chance to find the center of an object, and a size map, representing the width and height of an object at each location. CenterTrack [21] first adds a new input dimension, which corresponds to the heatmap of the object in the previous frame, and then it introduces a new branch to compute the object displacement between the current and the previous frame. The main benefit of this approach is that it does not require to be trained on video sequences, but it can be trained using static images with object detection ground truth. This allows the model to exploit large image datasets, like the CrowdHuman [23], and use the pre-trained CenterNet as a solid baseline for the training phase. The possibility of using object detection datasets represents the greatest strength of this model. On the other hand, the large amount of information to be provided as input and the computational cost of the inference step of such a big and complex neural network represents the main limitation of this algorithm in real-time applications.
Another well-known problem in detection-based trackers is the bias between the detection and the re-identification tasks. FairMOT [24] improved the performance of the tracking pipeline by implementing these two tasks within the same model using parallel branches. This approach solves the natural imbalance between the two tasks, making them ”fair” to each other. It also uses CenterNet [22] as a baseline, while introducing significant improvements in the model. For instance, the use of deformable convolutions [25] helped enhancing the receptive fields in the detection step, while the adoption of the Deep Layer Aggregation (DLA) [26], an enhanced version of the Feature Pyramid Networks (FPN) [27], improved the re-identification task significantly reducing the number of ID switches.
Transformers [28] were also used to learn complex tasks combining different steps of the object tracking pipeline. For instance, TransTrack [29] and Trackformer [30] are great examples of the use of such a technology. Both are based on the Detection Transformers (DETR) [31], which is one of the most prominent approaches for object detection. In particular, TransTrack [29] adds another transformer decoder to the DETR [31] architecture and introduces a matching head for producing tracking results. Trackformer [30] is an improvement of TransTrack [29] that uses just one transformer decoder, which directly produces the tracking queries. The set of queries obtained at time $t-1$ is then used in the current frame to match existing tracks or create new ones. However, transformers are complex models that require substantial computational resources to run in real-time, often not available in embedded devices. As also shown in the TransTrack paper, varying the number of decoders between 1 and 6, decreases the speed from 15 fps to 10 fps on a single V100 GPU, while the Trackformer executes at 7.4 fps.
II-B Separate trackers
Separate MOT algorithms track multiple objects independently by associating object detections in subsequent frames based on factors such as appearance, motion, and displacement. In separate trackers, each object is a standalone entity and its trajectory is determined without considering interactions or dependencies with other objects. In these algorithms, there is a clear separation between object detection, re-identification, and association logic. The evolution of deep neural networks occurred in the last decade had a significantly positive impact on tracking by detection algorithms.
DeepSORT [5] was the first solution able to demonstrate the advantages of deep learning in MOT. In particular, the use of deep neural networks allowed extracting more meaningful features from the tracked objects, leading to the adoption of more effective association metrics capable of reducing ID switches and mismatches, which represented the main weakness of SORT [4]. DeepSORT [5] was updated several times over the years to keep up with the evolution of increasingly powerful object detectors and re-identification networks.
To date, the most significant example of these updates is represented by StrongSORT [14], which introduced several improvements, like noise scale adaptive Kalman filter (NSA Kalman), exponential moving average weighting (EMA) of the embeddings, YOLOX [32] as object detection model, and BoT [33] as re-identification network. However, DeepSORT was known to be a very powerful and resource demanding framework and, despite the evolution of computing platforms, both DeepSORT and StrongSORT cannot provide sufficient real-time performance for cyber-physical systems with fast dynamics and limited computational resources, like rovers or drones.
The inference speed represents the pivotal point of BYTEtrack [17], which proposes a simple yet effective association metric only based on the intersection over union (IoU). Differently from the other state of the art solutions, BYTEtrack completely removes the re-identification network, thus saving a lot of computational resources. In fact, the only use of motion similarity allows the tracker to run at 30 fps on the MOT17 test set using a single V100 GPU. Note that high frame rates are also convenient because they lead to smaller changes in object positions between subsequent frames. On the other hand, the use of such a simple association metric does not work well in the presence of mid/long-term occlusions.
As a matter of fact, occlusions represents one of the most challenging situations that trackers have to deal with. If it is true that short-term occlusions can be managed effectively through a good motion model, (e.g., a linear Kalman filter and its extensions), long-term occlusions require the adoption of more advanced strategies.
Long-term occlusions are addressed by the observation-centric MOT, or OC-SORT [18], which improves the tracking robustness in complex scenarios where objects have non-linear motion and can interact with each other. This provides a better prediction capability in the presence of long-term occlusions, but the lack of a feature-based re-identificator represents its main limitation of this method. In fact, in situations where the occluded object changes its appearance significantly, the only motion model is not sufficient for recovering the track.
To address this problem, the same authors proposed Deep-OC-SORT [15], which introduces an adaptive appearance similarity-based association model upon OC-SORT [18], improving the tracking performance on the DanceTrack [34] benchmark by 6 HOTA [35] points. It also resulted to be the best model, according to the HOTA metrics, on both MOT17 [36] and MOT20 [16] benchmarks, clearly showing the importance of having a strong re-identification model. To date, state of the art re-identificator techniques are all based on deep neural networks, which by far outperform handcrafted or pure mathematical-based strategies. Their main drawback, however, is their computational cost.
This work proposes a new way to formulate the MOT problem more efficiently by splitting the Hungarian algorithm into multiple class-based matching tasks, which can then be parallelized, thus significantly reducing the overall computation time of the algorithm. Re-identification is used in the proposed tracking pipeline, due to its proven effectiveness, but the number of its invocations is reduced to save computation time thanks to a two-stage cascade matching. It worth observing that the proposed solution is general and can also be adopted to other state-of-the-art trackers to improve their inference time without affecting their accuracy.
III Proposed approach
As mentioned above, the proposed solution (named CORT) consists of splitting the Hungarian algorithm into a number of smaller class-based matching tasks, which can be parallelized to reduce the overall computation time of the algorithm. In order to better understand the proposed idea, the following section briefly recalls the Hungarian algorithm used in most separate trackers to optimally match bounding boxes detected at the current frame with bounding boxes predicted in the previous frame.
III-A The Hungarian match
The Khun-Munkres algorithm [37], also known as the Hungarian match, is a combinatorial optimizer used in separate trackers to find the optimal association between the bounding boxes retrieved by the object detection network at the current frame and those predicted by the tracker in the previous frame. The problem is formulated as a bipartite graph represented in a matrix form.
More specifically, given $p$ predictions and $d$ detections, let $c_{ij}$ be the cost for each (prediction, detection) pair computed according to a suitable metric (e.g., cosine similarity, Mahalanobis distance, or IoU). All costs can be represented as a $n\times n$ matrix $C$, where $n$ = max($p$, $d$), while the solution can be represented by a binary matrix $X$, with the same dimensions of $C$, where $x_{ij}=1$ means that the prediction $i$ best matches with the detection $j$.
Note that, if $p<d$, that is, if the predictions in the previous frame are less than the detections found in the current frame, then, for each row $i>p$, the cost $c_{ij}$ is set to a dummy value $V$ equal to the largest value computed for the elements in the previous rows, plus an arbitrary positive constant $k$, that is
$V=max\{c_{ij}|i=1,\ldots,p,j=1,\ldots,d\}+k.$ (1)
The exceeding detections will become new tracklets, if they persist for a given number of frames. Vice versa, if $p>d$, that is, if there are more predictions than detections, then, for each column $j>d$, the cost $c_{ij}$ is set to $V$ and the exceeding tracklets are maintained for a given number of frames (to manage temporary occlusions) and then deleted, if they do not match with any detection.
At each iteration of the tracking loop, a constrained optimization problem of the form reported in Equation (2) must be solved.
$\begin{cases}min\sum_{i=1}^{n}\sum_{j=1}^{n}c_{ij}x_{ij}\\ \sum_{j=1}^{n}x_{ij}=1\quad\forall i=1,\ldots,n\\ \sum_{i=1}^{n}x_{ij}=1\quad\forall j=1,\ldots,n\end{cases}$ (2)
The constraints of the problem impose that one and only one detection must be associated with exactly one prediction and vice versa. The computational complexity of the algorithm to find the optimal matching that minimizes the cost function is $O(n^{3})$.
III-B Concurrent MOT formulation
The Hungarian match algorithm is normally implemented as a monolithic approach that can hardly exploit the computational resources available in a multi-core platform, due to the intrinsic sequential nature of the operations involved in the process. To address such a problem, CORT proposes a parallel formulation of the Hungarian match in which the similarity matrix is divided into a number of smaller matrices, one for each object class. In this way, the solution of each subproblem is assigned to a different thread that can be executed in parallel with the other threads on a multi-core platform. Note that the number of possible object classes is established a priori and defines the maximum number of threads in which the solver can be split. The maximum performance gain is achieved when the number of threads is no larger than the number of available cores and the objects to be tracked are uniformly distributed among the various classes.
Consider, for instance, the typical situation in which each prediction has a corresponding detection to be associated with ($p=d$). In this case, all the objects are successfully tracked and there are no tracks to be deleted or created. Figure 2 shows how the data are organized in the traditional approach (left) and in the proposed solution (right).
Figure 2: Traditional approach for organizing the Hungarian matrix (left) and the proposed class-based formulation (right).
In the example illustrated in Figure 2, colors are associated with classes and it is clear that, when the objects in the scene belong to different classes, the number of values to be computed in the standard formulation is higher than the one required for the proposed formulation. Such a difference tends to increase with the number of objects to be tracked, hence the advantage of the class-based formulation tends to become more significant. A quantitative analysis of the resources required by the two approaches is presented in the next section.
Now, consider a situation in which the number of detections is higher than the number of predictions ($d>p$), because a new object entered the scene and requires the creation of a new tracklet. This is a very common situation that frequently occurs in a MOT task. For instance, assume that there are four predictions from the previous frame and five detections in the current frame, and assume that all the pre-existing tracks are successfully associated with one detection. This example is illustrated in Figure 3.
Figure 3: Differences in the two stages of the cascade matching for the standard monolithic approach (left) and the proposed class-based solution (right), when a new object enters the scene, so that $d>p$.
Adding a new element in the standard cost matrix involves creating a new row ($P_{5}$), filling all its elements with the dummy cost $V$, and then computing the similarity score between the new prediction ($D_{5}$) and all the pre-existing predictions $(P_{1},\ldots,P_{4})$.
A similar process is required in the dual case in which there are more predictions than detections ($p>d$), because a tracked object is occluded or exits the scene. In this case, a new column has to be added to the cost matrix and the exceeding prediction has to be compared against all the pre-existing detections to compute the corresponding similarity scores.
III-C Remarks on resource saving
The class-based implementation of the Hungarian matching algorithm proposed in CORT is able to save both computation time and memory, so achieving better time and space scalability. In particular, the following three factors concur to reduce computation times:
  1. 1.
    Splitting the matrix into class sub-matrices reduces the total number of elements, hence it reduces the number of similarity scores to be computed for each element.
  2. 2.
    In the class-based approach, each sub-matrix can be handled by a dedicated thread that can be executed in parallel with the others on a multi-core platform.
  3. 3.
    Since the complexity of the Hungarian algorithm is $O(n^{3})$, solving smaller sub-matrices amplifies the benefit in terms of overall execution time, even for a sequential execution.
Because of that, the execution time of the matching problem in the proposed solution is given by the time needed to solve the largest sub-problem, plus a small interference factor due to possible resource conflicts (i.e., cache, memory, bus) caused by the concurrent execution of multiple threads.
In the case considered in Figure 2, the solution of the standard Hungarian match has a time complexity proportional to $2166^{3}=$ 216 steps, while the proposed class-based solution only requires $243*2^{3}=$ 24 steps, which reduce to 8 steps, if the three matching sub-problems can run in parallel on three different cores.
In general, if $M$ is the number of possible classes, the maximum performance gain of the proposed approach is achieved when the $n$ objects to be tracked are uniformly distributed among the $M$ classes, so that the Hugariam matrix is split into $M$ sub-matrices, each of size $(n/M)\times(n/M)$. In such a best case, the memory used by CORT is $M$ times smaller than the monolithic case, while the time complexity reduces by a factor $M^{2}$. Also notice that the worst-case scenario for the proposed approach occurs when all the tracked objects belong to the same class. However, the worst-case complexity is not higher than the standard monolithic approach.
III-D Proposed tracking pipeline
The second contribution of this work is to propose a new cascade matching pipeline that uses the IoU at the first stage and the appearance model in the second stage, thus limiting the total number of operations and the execution of the cosine similarity matches.
As already proven in BYTEtrack [17], the use of the IoU match in the first stage is quite effective for high-confidence detections, since most of the objects do not normally require complex association metrics to be solved. Differently than BYTEtrack, in the proposed approach, the remaining tracks that are passed to the second stage are matched through the appearance model, using the cosine similarity as metric. In this way, the appearance matching is performed only for the remaining tracklets, thus reducing the overall workload.
Figure 4 shows the workflow of the two-stage cascade matching implemented in a generic thread at each iteration of the tracking loop. At the first stage, each cost matrix is populated only using the IoU association metric, where the cost computed for a generic (prediction, detection) pair $(P_{i},D_{j})$ is given by
$c_{ij}=1-IoU_{ij}.$ (3)
Figure 4: Workflow of the two-stage cascade matching proposed for the multi-class tracking algorithm.
In the proposed solution, the metric used for the localization matching is the Complete-Iou (CIoU) [38, 39], which improves the IoU by accounting for the distance between bounding boxes and their aspect ratio. Such a metric was proven to be more effective than IoU [40] and it has been used in other tracking algorithms, as OC-SORT and DeepOC-SORT [18, 15]. Note that, when using the CIoU, which varies between -1 and 1, the cost for localization matching is still defined as
$c_{ij}=1-CIoU_{ij},$ (4)
so that it varies from 0 to 2. In the first stage, the feature array for all the elements is also extracted to maintain a history of $h$ elements, updated for each track. In the second stage, the cosine similarity score is computed for each element in the history with respect to the detection that has to be matched, and the corresponding matching cost is computed as
$c_{ij}=1-\max_{k}\left(\frac{f_{ik}\cdot f_{j}}{\lVert f_{ik}\lVert\cdot\lVert f_{j}\lVert}\right),\forall k\in[0,h-1]$ (5)
where $f_{ik}$ is the $k$-th feature array in the history for track $i$, while $f_{j}$ is the feature array extracted from the current detection $j$. Resolving associations in the second stage based not only on the object position, but also using an appearance model, is crucial to increase the tracking performance and allows the tracker to recover from mid/long term occlusions.
IV Experiments
This section reports the results of three experiments carried out to validate the proposed approach. The first experiment aims at evaluating the time and memory scalability of CORT with respect to the traditional monolithic approach. The second experiment compares CORT with BYTEtrack in terms of both accuracy and execution time, whereas the third one compares the proposed solution with the standard approach in critical situations of unstable detection. All the results were obtained on a computing platform equipped with an i7 CPU and an Nvidia RTX-3060 Ti GPU.
Two different types of scenarios were used to validate the proposed solution: some experiments were executed on real-world videos, like those of the MOT challenge benchmark [16], while others were executed on synthetic videos generated using the CARLA simulator [41]. The use of CARLA was necessary to precisely control the various objects in the scene and test the performance of the proposed solution by varying the number of objects belonging to a specific class, as well as the total number of objects.
IV-A Evaluating time and memory scalability
The aim of this experiment is to show the effectiveness of the proposed solution for a different number of tracked objects in each class. In particular, the experiment started by generating an ideal scenario where the number of objects are equally distributed among the classes, and progressively changes it towards an unbalanced situation where the number of objects of a given class becomes more dominant.
Figure 1 shows the ideal scenario for CORT generated in the experiment, where six objects are present in the scene, equally distributed among three classes (car, bus, motorbike). In this case, the traditional Hungarian match formulation requires the creation of a $6\times 6$ cost matrix for a total of 36 entries representing the association scores between predictions and detections. Using the class-based approach, the problem requires the use of three $2\times 2$ cost matrices, for a total of 12 entries (4 + 4 + 4).
Table I reports the execution times required to populate the Hungarian matrix (or matrices, for multi-class approach) and solve the assignment problem(s). Each number reported in the table is the average over 50 executions. The two approaches have been executed in seven different scenarios, each identified as ($N_{1}$, $N_{2}$, $N_{3}$), where $N_{i}$ denotes the number of objects of class $C_{i}$ present in the scene ($C_{1}$ = ”car”, $C_{2}$ = ”bus”, $C_{3}$ = ”motorbike”).
It is worth observing that, in the best-case scenario (2, 2, 2), the time reduction reported for the proposed method is caused not only by the parallel execution of the threads in the corresponding cores of the processing platform, but also by the reduced size of the matrices, as explained in Section III. Although one could expect similar execution times for the three involved classes, this is not necessarily true. In fact, in CORT, the execution time depends both on the number of objects in the class and how many of them are forwarded to the second matching stage.
Scenario Proposed solution DeepSORT
Car Std Car Bus Std Bus Motorbike Std Motorbike Match time Match time Std
(2,2,2) 1.56 0.52 0.91 0.21 1.15 0.58 1.56 12.51 8.27
(3,2,1) 1.64 0.27 0.75 0.32 0.66 0.51 1.64 11.67 6.46
(1,1,4) 1.26 0.49 0.53 0.22 1.52 0.59 1.52 11.33 7.49
(6,0,0) 1.94 0.26 0 0 0 0 1.94 11.42 7.31
(0,0,6) 0 0 0 0 2.27 0.26 2.27 11.83 6.87
(3,3,3) 1.63 0.57 1.14 0.33 1.26 0.82 1.63 19.03 5.63
(4,4,4) 2.23 0.68 1.86 0.43 2.49 0.75 2.49 28.39 6.80
TABLE I: Execution times in milliseconds for processing the Hungarian match in the proposed multi-class parallel approach (CORT) and in the monolithic DeepSORT solution.
Under more general scenarios, when the number of objects in each class is unbalanced, the execution time of the multi-class approach is given by the largest class matching problem. Figure 5 shows a frame of the CARLA video sequence where there are three cars, two buses, and one motorbike.
Figure 5: A frame of the video recorded using the CARLA simulator where the subdivision by class is unbalanced.
The results of this test are also reported in Table I at the row labeled as (3, 2, 1). For the nature of the parallel solution, we can also say that increasing the number of objects in the other classes does not affect the timing performance of the algorithm significantly, as long as their number does not exceed the number of elements in the majority class and the new elements do not require intensive use of the second stage of the pipeline described in Section III. In contrast, in the monolithic approach, increasing the number of elements, regardless of their class, always increases the overall computation time of the algorithm. This is confirmed by the results reported in Table I; in fact, the execution time measured for the proposed multi-class approach for the case (3, 3, 3) is comparable with the time achieved for (3, 2, 1), whereas for the monolithic approach, the time for the case (3, 3, 3) results to be larger than the one for (3, 2, 1). This trend becomes even more evident in the (4, 4, 4) scenario, which includes 12 objects in the scene. Other two tests were carried out using different class distributions, as (6, 0, 0) and (0, 0, 6). Note that these scenarios represent the worst-case situation for the proposed approach, since all the elements belong to the same class. In this case, the resulting execution time of the multi-class method is reduced only by the different implementations of the cascade matching. In fact, for the (6, 0, 0) scenario, the proposed solution can solve the association problem using just the IoU matching for the majority of the frames, saving the time required by the second stage operations. This happens because cars are easier to detect with respect to motorbikes, leading to high confidence predictions from the Kalman filter. Vice versa, motorbikes, being smaller than cars and buses, are more likely to fail the first stage matching based on bounding box localization, so they are passed to the second stage, based on the appearance. This is why the execution times recorded for the (0, 0, 6) scenario are larger than those found for the (6, 0, 0) scenario, even though the number of objects is exactly the same. Finally, note that from the DeepSORT perspective, the execution times of the (6, 0, 0) and the (0, 0, 6) scenarios are comparable, since all the objects undergo the same computation, independently of their class.
IV-B Comparison with BYTEtrack
This experiment has been carried out to compare the execution time and the accuracy of CORT with the one of the best state of the art algorithm for real-time applications, that is BYTEtrack.
The main difference between the two tracking algorithms is the way they formulate the Hungarian match. In fact, BYTEtrack uses a decision threshold to split the detections into high and low confidence ones, and it never uses a neural network to perform a feature-based re-identification step. Differently, the proposed solution first divides the problem by classes and then performs the two-stage cascade matching explained in Section III. It is worth noticing that the choice of not using a neural network to get an appearance model exposes BYTEtrack to a higher number of id-switches when objects overlap to each other or in the presence of long-term occlusions.
Consider, for example, the scenario represented in the three frames reported in Figure 6.
Figure 6: An example of id-switch caused by the use of just the IoU metric, which can be avoided with a neural re-identification network and the cosine similarity.
In this case, because of the progressive overlap, the position of the two objects becomes rapidly imprecise due to the absence of a measure to correct the Kalman filter prediction. During this transient, the bus can still be matched, while the the two cars under occlusion are updated by dead reckoning. When the occluded objects becomes visible again, they reappear in a position that is far from the last detection. Since BYTEtrack only uses the IoU match, the chance of missing the recovery of the track is high, especially for small objects. In fact, if the motion of the hidden objects results to be highly non-linear during the occlusion, its predicted position is likely to have a negligible IoU with the detected one or a high IoU with the wrong track, possibly causing an id-switch. In contrast, the solution implemented in CORT can recover the track in the second stage of the cascade matching, because it exploits the features extracted by the re-identification network. Then, using the cosine similarity, it is able to get a more reliable association score, so recognizing a previously occluded object, rather than creating a new track. The situation described above is very common in multi-object tracking and the results reported in Table II confirm that the proposed solution improves the overall performance with respect to BYTEtrack by significantly reducing the number of id switches, with a small increase in the inference time.
Tracker FPS IDS MOTA MOTP
Proposed solution 17.8 3004 72.30 79.126
BYTEtrack 19.1 22989 62.79 80.329
TABLE II: A comparison of the BYTEtrack and the proposed class-based solution using different evaluation metrics.
For the sake of fairness, the results reported in Table II have been obtained by replacing the detection network of BYTEtrack with the one used in the proposed architecture, which is a YOLOv8 trained on the COCO dataset [42]. In this way, both trackers worked with the same set of detections. The achieved results also show that the re-identification stage slightly increases the execution time of the proposed tracker with respect to BYTEtrack. To evaluate such a difference, however, one should consider that BYTEtrack is optimized for tracking multiple objects belonging to a single class, while the proposed model is optimized for multi-classes tracking.
IV-C Behavior under unstable detections
This section examines the behavior of the standard and the multi-class approach in two critical situations in which the detection network could provide sporadic erroneous outputs in a frame sequence. To show the potential impact of such events on the performance of the tracker, consider a scenario similar to the one illustrated in Figure 7, where the red car appears in the scene at time $t_{a}$, but it is misclassified as a bus at time $t_{a}+2$. Also assume that all the other objects are correctly classified across the sequence. For the sake of clarity, only the detection responsible for the problem is shown in all the frames.
Figure 7: Evolution of the Hungarian matrices in the presence of unstable detections for the standard monolithic approach and the proposed multi-class solution.
To better understand the negative impact of such misclassification, Figure 7 also reports the evolution of the Hungarian matrices for both the standard monolithic approach and the proposed class-based solution, highlighting for each frame how this specific detection affects the size of the matrices and, hence, the number of operations required to fill them.
Note that, at time $t_{a}$, there are four predictions and five detections, due to the arrival of the new entity. As described in Section III-A, this requires the creation of a dummy row in the cost matrix to start the initialization phase of the new entity. Four of these five detections will be matched by the Hungarian algorithm with the predictions related to the pre-existing objects. The remaining one is used to create the data structure for the new track. In all trackers, the initialization phase requires a given amount of subsequent frames in which the new object must match with a detection. Such a transient is used to reject sporadic false detections and is also needed to initialize the covariance matrix of the Kalman Filter and adapt it to the motion of the new entity. If a new object is misclassified in such a transient, the track under initialization does not satisfy the requirements to be promoted as an active track. The result is that the object will be tracked with a larger delay, or its tracklet will never be activated, if the problem continues to occur. Note that, although this problem is common to all tracking algorithms, the multi-class solution is less affected in terms of computation with respect to the monolithic approach. In fact, in the multi-class match, a class change increases the dimension of two small matrices by one, whereas in the monolithic approach, the dimension of the Hungarian matrix is always increased by two.
Consider now a slightly different scenario in which all the objects in the scene are correctly tracked. Suppose that, for a given frame, the following two events occur for two close objects at the same time:
  • Missed detection: one of the object is not detected in the frame.
  • Wrong classification: the detection having an IoU value greater than zero with the two predictions changes its class.
Under this assumption, we have two predictions of different classes, but only one detection that can be used for updating the Kalman filter. If the detection contended between the two predictions changes its class faster than the number of frames tolerated by the tracker to remove a track, then a track that should have been eliminated can persist longer than needed. Such a persistent track also causes additional operations to be performed that could have been avoided. Notice that, also in this scenario, the proposed class-based solution behaves better than the standard monolithic approach. In fact, since the standard method only uses one matrix, maintaining a track active is more expensive in terms of operations, and this becomes worse as the number of objects in the scene increases. Vice versa, the multi-class approach requires fewer computations because the class matrices are always smaller than the one in the monolithic solution.
A simple solution to attenuate the two problems described in this section is to increase the acceptance threshold of the detection network. Raising the value of such a threshold ensures that only reliable detections are passed to the tracker to compute the costs for the assignment problem. At the same time, raising the threshold increases the number of false negatives. However, this is not a major issue, because the Kalman filter can efficiently compensate for some missing detections. This is the reason why in Table II the value for the precision (MOTP) is lower than that of the BYTEtrack, while the accuracy (MOTA) is higher because of the reduced number of id-switches.
V Conclusions
This work presented CORT, an efficient method to enhance the timing performance of modern real-time tracking algorithms. Such an enhancement is obtained by splitting the Hungarian matching algorithm into smaller association problems, carried out per class. Considering that the matching algorithm has an $O(n^{3})$ time complexity, where $n$ is the number of elements in the cost matrix, and that each sub-matrix can be solved in parallel on different cores, the proposed approach significantly improves the performance of state of the art tracking by detection systems. As highlighted in Section IV, as the number of objects to be tracked increases, the proposed parallel solution becomes more and more efficient than the monolithic one. This is also confirmed by the experimental results reported in Table I, showing that the gap between the execution time of the two methods significantly increases for tracking scenarios with a larger number of objects, highlighting the better scalability of the proposed solution.
As a future work, we plan to accelerate on FPGA some vectorizable operations of the proposed pipeline, as the computation of the association scores. Another optimization that can be further investigated is the dynamic tuning of the hyperparameters of the algorithm. In fact, as reported in Section IV-C, changing these values has a direct impact on the behavior of the tracker. A runtime adaptation would be desirable to maximize the amount of information extracted from each frame while avoiding the problems reported in Section IV.
VI Acknowledgments
This work was partially supported by the Department of Excellence of Robotics and AI of the Sant’Anna School of Advanced Studies.